# LaCour!: Enabling Research on Argumentation in Hearings of the European Court of Human Rights

### v1

@@ Abstract
Why does an argument end up in the final court decision? Was it deliberated or questioned during the oral hearings? Was there something in the hearings that triggered a particular judge to write a dissenting opinion? Despite the availability of the final judgments of the European Court of Human Rights (ECHR), none of these legal research questions can currently be answered as the ECHR’s multilingual oral hearings are not transcribed, structured, or speaker-attributed. We address this fundamental gap by presenting LaCour!, the first corpus of textual oral arguments of the ECHR, consisting of 154 full hearings (2.1 million tokens from over 267 hours of video footage) in English, French, and other court languages, each linked to the corresponding final judgment documents. In addition to the transcribed and partially manually corrected text from the video, we provide sentence-level timestamps and manually annotated role and language labels. We also showcase LaCour! in a set of preliminary experiments that explore the interplay between questions and dissenting opinions. Apart from the use cases in legal NLP, we hope that law students or other interested parties will also use LaCour! as a learning resource, as it is freely available in various formats at https://huggingface.co/datasets/TrustHLT/LaCour.
@@ 1 Introduction
In order to better understand the legal argument mechanics, that is which arguments of the parties were presented, discussed, or questioned, and thus might have influenced the case outcome, we must take the oral hearings into account. We witness that the availability of oral hearing transcripts of the U.S. Supreme Court enables further legal research (Ashley et al, 2007). However, empirical research into the interplay of arguments at the court hearings and the final judgments has been so far impossible for the ECHR, as there are no hearing transcripts available.
@@ 2 Related work
Easily accessible and extensive collections of all court material naturally play an important role for legal scholars. But the transcripts alone are also of great importance for empirical research, as they enable such research for NLP in the first place. Some approaches have used the transcripts of the oral arguments to gain insights into the argument structure for the purpose of tutoring (Ashley et al, 2007), others have tried stance detection (Bergam et al, 2022) or machine learning-based approaches to study the judges’ attitudes (Dickinson, 2018).
@@ 2.2 Hearings of the ECHR
On the European side of legal analysis in the context of NLP, a popular choice of investigation is the ECHR. [^7] https://www.echr.coe.int/ This court is a frequently used source for analysis, not only because of the availability of its documents, but also because of its importance and reputation as a supranational body. It has been in existence since 1953 and deals with applications brought by individuals or groups, in rare cases even states, against states that are contractors to the European Convention on Human Rights (“The Convention”). The Convention contains several articles [^8] https://www.echr.coe.int/european-convention-on-human-rights concerning different aspects of human rights and applications can refer to multiple of those articles. The court receives over 50,000 applications every year which are proceeded according to importance and urgency. For roughly 30 Chamber or Grand Chamber cases each year, the court holds a public hearing which is recorded.999more information on how an application is processed can be found in https://www.echr.coe.int/documents/d/echr/your_application_eng
@@
Another popular use of ECHR material is Legal Judgment Prediction (LJP), most notably investigated by Aletras et al (2016). Besides classifying cases in a binary manner for violation of a specific Article, they also introduce a publicly available corpus with 584 cases. All cases in this dataset are in English, relate to Articles 3, 6 and 8 of the Convention and are balanced in regards to violation and non-violation. Chalkidis et al (2019) collect 11.5k cases in a similar manner and use them for binary classification, multi-label classification and case importance detection with the help of neural models. Medvedeva et al (2020) use a similarly sized portion of ECHR cases and experiment with only using information available before the final judgment to simulate actual prediction rather than classification. In an additional experiment, they also investigate a connection of a specific judge being part of the Chamber and the decision of the case. The latter experiments even achieved a surprisingly high correctness for certain Articles of the Convention. However, prediction for ECHR cases still suffers from the issue that argumentative text is only available after the judgment has been passed and anything using the judgment documents is already biased towards the outcome. Moreover, as Villata et al (2022) pointed out in their review of (Medvedeva et al, 2020), classifying legal cases is a different and more complex matter than a simple binary classification.
Finally, ECHR documents are also a popular source used in large dataset collections or benchmarks (Chalkidis et al, 2022; Henderson et al, 2022), which made them a standard and go-to option for many works in legal NLP. And due to the amount of available text, they have also been used as training material for domain specific language models, like LEGAL-BERT (Chalkidis et al, 2020) which have opened up new possibilities for handling legal texts.
However, both legal argument mining and legal judgment prediction focus on the written documents provided by the ECHR through HUDOC. Because there are no transcripts of the public oral arguments in hearings in the ECHR available, to the best of our knowledge, no such research has been conducted in this area and although the hearings are publicly available, they have not been scrutinized in NLP research to date. The hearings also present the gap in the current representation of legal proceedings in the ECHR. Although held only on a proportionally small number of cases, oral hearings can provide insight into the process that takes place between an application and the final judgment.
@@ 3.1 Hearing videos
As our source material, we collected all videos of hearings - called webcasts - that are available on the website [^11] https://www.echr.coe.int/webcasts-of-hearings. At the time we started the collection, the available hearings ranged from 2007 to 2022 with 3 different audio tracks: the original version without interpretation, an English interpretation and a French interpretation. There are no official transcripts of the hearings and the videos do not contain any subtitles. The only available material are live recorded videos with the live recorded interpretations as an optional audio track.
@@ 3.2 Generating transcripts
The very first step in the processing pipeline, illustrated in Figure 2, is the speaker diarization, which is applied to obtain a schedule of who is speaking at which time. The diarization also takes care of overlapping speech that occurs in rare cases. For this task, we ran pyannote [^12] https://huggingface.co/pyannote/speaker-diarization (Bredin et al, 2020) which returns a fine-grained schedule of speech segments and a supposed speaker identification number. We clustered the output of pyannote by speaker change into larger chunks to create a “speaker schedule” that resembles the order of speeches in the hearing.
@@
To account for the different languages spoken, we originally wanted to use automatic language detection on those segments. However, initial trials with several wav2vec2 models to detect the spoken language did not yield sufficiently reliable accuracy. And even though the field of speech and language recognition has seen large improvements with models like Whisper (Radford et al, 2023), we still encountered a number of issues with the language detection in Whisper as well. One major problem is the fact that participants of the hearings, especially the judges, tend to speak both English and French and not seldom use both languages in a single sentence. A common example of this is “Merci beaucoup, Madame. Judge [Name], please.” The automatic language detection for utterances like this failed frequently. The second most common problem is related to the language proficiency of the speaker. With English and French being the official languages, many speakers before the Court do not speak in their native language. And although recognizing the spoken language in these cases might not be an issue for a human, the automatic language detection often picks up the native language of the speaker, rather than the actual spoken language. We suspect the detection model might be biased towards native speakers.
@@
To circumvent this unreliable detection, we manually identified the languages for each section in our coarse-grained speaker schedule. We hired two research assistants to listen to parts of the speaker segments in order to identify the spoken language as well as the affiliation of the speaker, which would later be needed to distinguish a party’s address from the judges’ questions. Any conflicting identification by the annotators was discussed until a correct tag was agreed on. Additionally, this manual step allowed us to correct small diarization errors of the framework. The resulting speaker schedules contain segments which are marked with timestamps, as well as the spoken language (ISO 639-1) and the role of the speaker. For the roles, we decided to distinguish between the parties appearing in court. These are Announcer, Applicant, Government, Judge President (“JudgeP”), Judge and Third Party. If there are multiple parties as for example multiple applicants in a case, the tags do not distinguish between them. In the case “Savickis and Others v. Latvia (no. 49270/11)” for example, the representatives for “Savickis” and the representatives for all other applicants would be tagged simply as “Applicant”.
@@
The central component of the pipeline is the transcription itself. Our method for generating a transcript involves using Whisper (Radford et al, 2023) on a single entry of a speaker schedule, passing the previously identified language as a parameter instead of using the model’s language detection. For long segments of speech this produces a single large block of text. To ensure better alignment between timestamps and transcript, we switched to using a sentence segmenting version of Whisper [^13] https://github.com/jianfch/stable-ts to generate more accurately aligned sentence segments of the transcript. The resulting transcript is divided into speaker segments which are additionally divided into fine-grained sentence segments. One additional round of manual correction was necessary to properly separate instances of speaker overlap, as the model generates inaccurate transcripts when two people are speaking simultaneously. Segments with potential overlap were tagged automatically and evaluated by our student research assistants to resolve these conflicts.
@@
Even though this transcript version is sufficient for understanding the hearings, we decided to further refine it by adding the names of the judges and correcting the transcript. To obtain the names, we collected lists of hearing participants from linked documents in the HUDOC databse titled “Forthcoming Hearing” and verified that the participants listed in them also appear in the “Judgement” document. This step was necessary to spell the names of the representatives correctly, particularly when they are mentioned in the transcript. The names of the judges could be used later to link judges’ opinions in the final judgments with questions asked by them. Although the produced transcripts are surprisingly accurate, legal terminology and named entities are frequently transcribed incorrectly, e.g. special terminology such as “tax arrears” is often transcribed as “tax errors”.
We hired two student research assistants proficient in English and French and tasked them with manually correcting the parts spoken by the judges and the judge president. They reported an average effort of 1,5 hours per hearing and corrected the parts of the entire corpus in approximately 300 hours. Due to limited resources, speeches by other speaker roles (“Applicant”, “Government”, “Third Party”) and in other languages remain uncorrected. Our completed transcript corpus comprises tagged speech segments with tagged sentence-aligned transcriptions, which are manually corrected for the labels “Judge” and “Judge President”, but uncorrected for all other labels. Figure 3 displays an example of a transcript. Obtaining the final partially corrected transcript from the video version required the work of 4 student research assistants and took about 10 months. We call our new corpus LaCour! in honor of the catchy announcement at the beginning of each hearing.
@@ 3.3 Linked documents
We collected documents from the HUDOC database that are associated with any application number that was attached to a hearing webcast. Due to different document types, there are multiple documents associated with a hearing. Some manual adjustments to the matching were required, because the official database occasionally links to an incorrect document. We exclusively retrieved English documents, but it is also possible to find the link to a French document, should it be available. During the process of acquiring the matching judgment files, we noticed some of the applications were deemed inadmissible or were withdrawn by the applicant after the hearing. Because having a hearing transcript without a conclusion in form of a judgment appeared insufficient, we excluded all hearings without a judgment file. The list of excluded hearings can be seen in Appendix A along with the reason for exclusion.
All judgment files are provided with their respective case-id and application number, the latter of which serves as a link to the hearing and webcast-id. A webcast only has a single id, while there can be more than one application number connected to a hearing or a case. It should also be noted that there is overlap [^14] We found an overlap based on the item-id for 190 out of 469 documents of these documents to the corpus by Chalkidis et al (2019), but unfortunately there is no overlap to the documents annotated by Habernal et al (2023). This is primarily because LaCour! mostly contains documents with an importance score of 1, meaning “key cases” and only from a limited date range. The case-ids and application numbers can also serve as a key to identify and collect more material, because it is the same identification as used by the court.
@@
Our transcriptions are not meant to reflect verbatim what is said in the audio, but to aid the understanding of the hearing in text form. We therefore do not correct the text to reflect stuttering, sentences that have been slightly corrected during speaking, filler words or incorrect grammar. But because Whisper relies on an internal language model, the generated sentences are sometimes grammatically correct, even though the actual spoken sentence is not. This is especially true for the parts we did not manually correct, which concerns everything spoken by the roles “Applicant”, “Government” and “Third Party”. Manually corrected and modified parts are sections spoken by “Judge” and “Judge President”, as well as all language tags. The remainder of the text has been left untouched and is a direct result of Whisper. Several format and completeness checks were additionally run on the files to ascertain that no tags are incomplete or missing and the judges’ names match the ones listed as attendants.
@@
In order to enable easy dissemination and also provide an easily readable version of a transcript, we provide LaCour! in different file formats. For comfortable reading, we refer to the concatenated version [^15] https://github.com/trusthlt/lacour-corpus or our online preview page. [^16] https://www.trusthlt.org/lacour/ The sentence aligned version is available in xml and text format on huggingface. [^17] https://huggingface.co/datasets/TrustHLT/LaCour At this point we would like to emphasize that the transcripts are unofficial interpretations of the hearing webcasts. The webcasts in conjunction with the provided documents are the only official presentation of the proceedings.

### v4

@@ Abstract
Why does an argument end up in the final court decision? Was it deliberated or questioned during the oral hearings? Was there something in the hearings that triggered a particular judge to write a dissenting opinion? Despite the availability of the final judgments of the European Court of Human Rights (ECHR), none of these legal research questions can currently be answered as the ECHR’s multilingual oral hearings are not transcribed, structured, or speaker-attributed. We address this fundamental gap by presenting LaCour!, the first corpus of textual oral arguments of the ECHR, consisting of 154 full hearings (2.1 million tokens from over 267 hours of video footage) in English, French, and other court languages, each linked to the corresponding final judgment documents. In addition to the transcribed and partially manually corrected text from the video, we provide sentence-level timestamps and manually annotated role and language labels. We also showcase LaCour! in a set of experiments that explore the interplay between questions and dissenting opinions. Apart from the use cases in legal NLP, we hope that law students or other interested parties will also use LaCour! as a learning resource, as it is freely available in various formats at https://huggingface.co/datasets/TrustHLT/LaCour.
@@ 1 Introduction
In order to better understand the legal argument mechanics, that is which arguments of the parties were presented, discussed, or questioned, and thus might have influenced the case outcome, we must take the oral hearings into account. We witness that the availability of oral hearing transcripts of the U.S. Supreme Court enables further legal research (Ashley et al, 2007; Fang et al, 2023b, a). However, empirical research into the interplay of arguments at the court hearings and the final judgments has been so far impossible for the ECHR, as there are no hearing transcripts available.
@@ 2 Related work
Easily accessible and extensive collections of all court material naturally play an important role for legal scholars. But the transcripts alone are also of great importance for empirical research, as they enable such research for NLP in the first place. Some approaches have used the transcripts of the oral arguments to gain insights into the argument structure for the purpose of tutoring (Ashley et al, 2007), others have tried stance detection (Bergam et al, 2022) or machine learning-based approaches to study the judges’ attitudes (Dickinson, 2018). More recently Fang et al (2023b) compiled a large collection of court documents of the U.S. Supreme Court, called Super-SCOTUS, which specifically highlights the role of hearing transcripts. In the work by Fang et al (2023a), Super-SCOTUS enables the study of the role and influence of partisanship in oral hearings.
@@ 2.2 Hearings of the ECHR
On the European side of legal analysis in the context of NLP, a popular choice of investigation is the ECHR. [^7] https://www.echr.coe.int/ This court is a frequently used source for analysis, not only because of the availability of its documents, but also because of its importance and reputation as a supranational body. It has been in existence since 1953 and deals with applications brought by individuals or groups, in rare cases even states, against states that are contractors to the European Convention on Human Rights (“The Convention”). The Convention contains several articles [^8] https://www.echr.coe.int/european-convention-on-human-rights concerning different aspects of human rights and applications can refer to multiple of those articles. The court receives over 50,000 applications every year which are processed according to importance and urgency. For roughly 30 Chamber or Grand Chamber cases each year, the court holds a public hearing which is recorded.999more information on how an application is processed can be found in https://www.echr.coe.int/documents/d/echr/your_application_eng
@@
Another popular use of ECHR material is legal judgment prediction, most notably investigated by Aletras et al (2016). Besides classifying cases in a binary manner for violation of a specific Article, they also introduce a publicly available corpus with 584 cases. All cases in this dataset are in English, relate to Articles 3, 6 and 8 of the Convention and are balanced in regards to violation and non-violation. Chalkidis et al (2019) collect 11.5k cases in a similar manner and use them for binary classification, multi-label classification and case importance detection with the help of neural models. Medvedeva et al (2020) use a similarly sized portion of ECHR cases and experiment with only using information available before the final judgment to simulate actual prediction rather than classification. In an additional experiment, they also investigate a connection of a specific judge being part of the Chamber and the decision of the case. The latter experiments even achieved a surprisingly high correctness for certain Articles of the Convention. However, prediction for ECHR cases still suffers from the issue that argumentative text is only available after the judgment has been passed and anything using the judgment documents is already biased towards the outcome. Moreover, as Villata et al (2022) pointed out in their review of (Medvedeva et al, 2020), classifying legal cases is a different and more complex matter than a simple binary classification.
Finally, ECHR documents are also a popular source used in large dataset collections or benchmarks (Chalkidis et al, 2022; Henderson et al, 2022; Guha et al, 2023), which made them a standard and go-to option for many works in legal NLP. And due to the amount of available text, they have also been used as training material for domain specific language models, like LEGAL-BERT (Chalkidis et al, 2020) which have opened up new possibilities for handling legal texts.
A case in the ECHR can be evaluated on more than the final decision. Factors outside of the content of a decision, such as voting behavior, opinions and circumstances which potentially influence a judge’s decision have also been studied empirically using decisions and opinions taken from HUDOC (Voeten, 2008).
Still, NLP-based techniques such as legal argument mining and legal judgment prediction focus on the written documents provided by the ECHR through HUDOC. Because there are no transcripts of the public oral arguments in hearings in the ECHR available, to the best of our knowledge, no such research has been conducted in this area and although the hearings are publicly available, they have not been scrutinized in NLP research to date. The hearings also present a data gap in the current representation of legal proceedings in the ECHR. Although held only on a proportionally small number of cases, oral hearings can provide insight into the process that takes place between an application and the final judgment.
@@ 3.1 Hearing videos
As our source material, we collected all videos of hearings - called webcasts - that are available on the website. [^11] https://www.echr.coe.int/webcasts-of-hearings At the time we started the collection, the available hearings ranged from 2007 to 2022 with 3 different audio tracks: the original version without interpretation, an English interpretation and a French interpretation. There are no official transcripts of the hearings and the videos do not contain any subtitles. The only available material are live recorded videos with the live recorded interpretations as an optional audio track.
@@ 3.2 Generating transcripts
Getting from a pure video to a structured and annotated transcript required the design and implementation of a processing pipeline. The very first step in said pipeline, illustrated in Figure 2, is the speaker diarization, which is applied to obtain a schedule of who is speaking at which time. The diarization also takes care of overlapping speech that occurs in rare cases. For this task, we ran pyannote [^12] https://huggingface.co/pyannote/speaker-diarization (Bredin et al, 2020) which returns a fine-grained schedule of speech segments and a supposed speaker identification number. As a next step, we clustered the output by a change in speaker into larger chunks to create a “speaker schedule” that resembles the order of speeches in the hearing.
@@
A particularly hard step in the pipeline was the language detection. We originally wanted to use automatic language detection to tag each segment of the speaker schedule. However, initial trials on selected samples with several wav2vec2 (Baevski et al, 2020) models [^13] We experimented with the huggingface models facebook/wav2vec2-conformer-rel-pos-large-960h-ft, facebook/wav2vec2-base-10-voxpopuli-ft-en, facebook/hubert-large-ls960-ft to detect the spoken language did not yield sufficiently reliable results. Even the much more powerful Whisper (Radford et al, 2023) with an in-built language detection in the pipeline had issues with many instances that we tested. One major problem is the fact that participants of the hearings, especially the judges, tend to speak both English and French and not seldom use both languages in a single sentence. A common example of this is “Merci beaucoup, Madame. Judge [Name], please.” The second most common problem is related to the language proficiency of the speaker. With only English and French being the official languages, many speakers before the Court do not speak in their native language. For a human, this is not a challenging task, automatic language identification models on the other hand had serious problems and frequently picked up the native language of the speaker, rather than the actual spoken language. As an example, French speeches by Italian native speakers were classified as Italian. We suspect that the detection models are biased towards native speakers, which could be caused by the data they are trained on, which is mostly clean speeches spoken by native speakers.
@@
We could not find a reasonable way to automatically label the language, without having to check every instance. We therefore decided to manually identify the languages for each section in our coarse-grained speaker schedule. To achieve this, we hired two research assistants who were trained to listen to parts of the segmented videos in order to identify the spoken language. They validated the correctness of the timestamps and adjusted them if necessary by directly modifying the schedule file. Simultaneously, they tagged the affiliation of the speaker, which would later be needed to distinguish a party’s address from the judges’ questions. The correct role could be inferred from nameplates in the video and the context of the president’s announced order. Any conflicting identification by the annotators was discussed until a correct tag was agreed on. This manual step also allowed us to correct small diarization errors. The resulting speaker schedules contain segments which are marked with timestamps, as well as the spoken language in ISO 639-1 format and the role of the speakers. For the roles, we decided to distinguish between the parties appearing in court. These are ‘Announcer’, ‘Applicant’, ‘Government’, ‘JudgeP’ (judge president), ‘Judge’ and ‘Third Party’. In cases in which there are multiple parties, as for example multiple applicants, the tags do not distinguish between them. In the case “Savickis and Others v. Latvia (no. 49270/11)” for example, the representatives for “Savickis” and the representatives for all other applicants would be tagged as “Applicant”.
@@
This tagged structure is sufficient, but also necessary for the central component of the pipeline: the transcription. Our method for generating a transcript is based on Whisper [^14] https://huggingface.co/openai/whisper-large-v2 (Radford et al, 2023) on a single entry of a speaker schedule, passing the previously identified language as a parameter. Passing the language tag circumvents the in-built language detection and ensures that that the output text is in the correct language. Inputting the typically long segments of speech results in a single large block of text. To ensure better alignment between timestamps and transcripts, we used stable-ts [^15] https://github.com/jianfch/stable-ts which generates more accurately aligned sentence segments of the transcript. As a result, the transcript is not just divided into speaker segments, but also into fine-grained sentence segments. A last manual correction step was necessary to properly separate instances of speaker overlap, as the model generates inaccurate transcripts when two speakers talk at the same time. We identified segments with potential overlap automatically using the timestamps and the annotators were tasked to separate and resolve these conflicts.
@@
Even though this transcript version is sufficient for understanding the hearings, we decided to further refine it by adding the names of the judges and correcting the transcript. To obtain the names, we gathered lists of hearing participants from linked documents in the HUDOC database titled “Forthcoming Hearing” and verified that the listed participants also appear in the “Judgement” document. Judges and names in general are often misspelled in the transcript and we used the listed participants to correct the spelling. This correction enabled us to later link judges’ opinions in the final judgments with questions they asked during the hearing.
We hired two research assistants with a high proficiency in English and French to manually correct all parts spoken by the judges or the judge president in the automatically generated transcripts. The research assistants reported an overall good quality and little corrections to be made. Legal terminology and named entities were occasionally inaccurately transcribed. For example, domain specific terminology such as “tax arrears” was transcribed as “tax errors”. Corrections took them approximately 1.5 hours per hearing on average, resulting in 300 hours invested for all hearings. Due to limited resources and issues in finding a suitable crowdsourcing opportunity, speeches by other speaker roles (“Applicant”, “Government”, “Third Party”) and in other languages besides English and French remain uncorrected. In parallel, we instructed the research assistants to add the names of the judges to the corrected segments. They were able to chose from the previously compiled list of attendants to make sure that there are no differences in spelling.
Finally, our completed transcript corpus comprises tagged speech segments with tagged sentence-aligned transcriptions, which are manually corrected for the labels “Judge” and “Judge President”, but uncorrected for all other labels. Figure 3 shows an excerpt of one transcript. Obtaining the final partially corrected transcript from the video version required the work of 4 student research assistants and took around 10 months. We call our new corpus LaCour! in honor of the traditional and catchy announcement at the beginning of each hearing.
@@ 3.3 Linked documents
We collected documents from the HUDOC database that are associated with any application number that was attached to a hearing webcast. Due to different document types, there are multiple documents associated with a hearing. Some manual adjustments to the matching were required, because the official database occasionally links to an incorrect document. We exclusively retrieved English documents, but it is also possible to find the link to a French document, should it be available. During the process of acquiring the matching judgment files, we noticed some of the applications were deemed inadmissible or were withdrawn by the applicant after the hearing. Because having a hearing transcript without a conclusion in form of a judgment appeared insufficient, we excluded all hearings without a judgment file. The list of excluded hearings can be seen in Appendix 2 along with the reason for exclusion.
All judgment files are provided with their respective case-id and application number, the latter of which serves as a link to the hearing and webcast-id. A webcast only has a single id, while there can be more than one application number connected to a hearing or a case. It should also be noted that there is overlap [^16] We found an overlap based on the item-id for 190 out of 469 documents of these documents to the corpus by Chalkidis et al (2019), but unfortunately there is no overlap to the documents annotated by Habernal et al (2023). This is primarily because LaCour! only contains documents with an importance score of 1, which are key cases and only from a limited date range. The case-ids and application numbers can also serve as a key to identify and collect more material, because it is the same identification as used by the court.
@@
Our transcriptions are not meant to reflect verbatim what is said in the audio, but to aid the understanding of the hearing in text form. We therefore instructed our annotators to not correct the text to reflect stuttering, sentences that have been slightly corrected during speaking, filler words or incorrect grammar. But because Whisper relies on an internal language model, the generated sentences are sometimes grammatically correct, even though the actual spoken sentence is not. This is especially true for the parts we did not manually correct, which concerns everything spoken by the roles “Applicant”, “Government” and “Third Party”. Manually corrected and modified parts are sections spoken by “Judge” and “Judge President”, as well as all language tags. The remainder of the text has been left untouched and is a direct result of Whisper. Several format and completeness checks were additionally run on the files to ascertain that no tags are incomplete or missing and the judges’ names match the ones listed as attendants. On top of that, we automatically checked all transcripts for unusual characters and repeated sentences resulting from looping behavior in the Whisper generative language model.
@@
Easy dissemination and readability has a high priority for us, so we provide LaCour! in different file formats. For comfortable reading, we refer to the concatenated version [^17] https://github.com/trusthlt/lacour-corpus or our online preview page. [^18] https://www.trusthlt.org/lacour/ The sentence aligned version is available in xml or text format on huggingface. [^19] https://huggingface.co/datasets/TrustHLT/LaCour At this point we would like to emphasize that the transcripts are unofficial interpretations of the hearing webcasts. The webcasts in conjunction with the provided documents are the only official presentation of the proceedings.